OpenStudy (anonymous):

what is the solution to the equation below? 5/4x - 5/2 = 2

OpenStudy (anonymous):

add 5/2 to both sides 5/4x = 5/2 +2 Simplify the right-hand side of the equation 4/4x = 9/2 5 · 2=9 · 4x 9 · 4x=5 · 2 multipy 5 by 2 to get 10 9 · 4x=10 multipy 9 by 4x to get 36 36x/36 =10/36 x= 10/36 simplfy divide by 2 X= 5/18

OpenStudy (anonymous):

they have 18/5 but not 5/18

OpenStudy (anonymous):

maybe X=10/36